A standard six-sided die is rolled 12 times. What is the standard deviation of the number of times a 2, 3, or 4 will be rolled? Round your answer to two decimal places.
Answer #1

This will be a binomial distribution with parameters:

n = 12, p = P(2, 3 or 4) = 3/6 = 0.5

Hence,

Standard deviation = = 1.73